OpenStudy (anonymous):

Write an equation of the line described in slope intercept form that passes through -7, -2 and is perpendicular to Y = -1/2X + 4. Can someone help me through this?

OpenStudy (anonymous):

Perpendicular slope to the given line--opposite reciprocal. so it would be 2. Now use y=mx+b

OpenStudy (dan815):

is that ellen in your picture

OpenStudy (anonymous):

y = 2x + b (-7,-2) -2 = 2(-7)+b -2 = -14+b 12 = b now you have m=2 and b = 12 y = 2x+12

OpenStudy (anonymous):

No...not Ellen. It's really me. People ask me that every now and then.

OpenStudy (anonymous):

You are an amazing help!!! Ty so much.

OpenStudy (anonymous):

You're welcome